OpenStudy (anonymous):

4(x-6)-(3x+2)=-5 step by step plz

OpenStudy (anonymous):

4*x - 4*6 - 3x - 2 = -5 4x - 24 - 3x - 2 = -5 x - 26 = -5 x = -5 + 26 x = 21

OpenStudy (aroub):

Distribute: 4x-24-3x-2=-5 put all the like terms on a side : 4x-3x=24+2-5 add like terms: x=21